WARNING: This web is not used anymore. Please use PDBService.MinuteS19June08 instead!
 

Minutes 19 June 08

  • Phone:
    • Gridka: Andreas
    • RAL: Carmine
    • NDGF: Olli
    • BNL: Carlos
    • TRIUMF: Andrew
    • SARA: Alexander

  • CERN:
    • 3D: Dirk, Maria, Luca, Eva
    • ATLAS: Gancho, Florbela

  • Site status:
    • Tier0:
      • Downstream databases upgraded to 10.2.0.4
      • Mail to the 3d list – parameter related to Streams to be included during Tier1s interventions in order to upgrade to 10.2.0.4
      • Migration of the ATLAS database at PIC using transportable tablespaces for the Streams resynchronization
      • Old databases still maintained in the Streams setup. Eva will contact Luis so we can schedule their removal
      • d3r RAC database being reinstalled with 10.2.0.4 64 bit
    • SARA (Alexander): nothing to report. 30 June, maintenance OPCORN network, downtime 2 hours. Upgrade to 10.2.0.4 in the next couple of weeks (to be scheduled)
    • RAL (Carmine): nothing to report. Upgrade 10.2.0.4 of the test machines. CPU patch to be applied in production next week. Upgrade to 10.2.0.4 intervention date to be decided (probably before end July). Plan to add new node to ATLAS database. Maria suggested to be added before end July
    • Triumf (Andrew): Upgrade to 10.2.0.4 before July 11th (date still to be decided)
    • BNL (Carlos): Upgrade to 10.2.0.4 already scheduled for 23rd June
    • NDGF (Olli): Upgrade to 10.2.0.4 in the next 2 weeks (before Olli goes on holidays)
    • Gridka (Andrea): Upgrade to 10.2.0.4 scheduled for July 19th. New person working on 3d project at Grika
  • Please check your support mail list in the wiki before going on holidays!!!!

  • ATLAS:
    • PVSS replication: From Online to Offline: try 1 account during 1 month.
    • Online database: patching has been delayed until news about this database (probably in the the next weeks)
    • Grant select on “view” bug on Streams is not fixed – this can be a problem with the PVSS replication. We have to think about a solution/workaround. Eva will try to increase the priority of the SR with Oracle support
    • Muon centers: joined to Streams replication to INTR. We do not have responsible persons so we have to wait to move to production
    • AMI: preparation of source database ongoing

*Next meeting in 2 weeks

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2008-06-26 - EvaDafonte
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    PSSGroup All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback